OpenStudy (anonymous):

What is the value of x? A. 5 B. 7 C. 10 D. 11 If you don't like giving direct answers, then please at least tell me how to get it so I can solve it myself.

OpenStudy (anonymous):

The fraction is x over 12 and equals 55 over 60.

OpenStudy (alexandervonhumboldt2):

x/12=55/60 x=55/60*12

OpenStudy (alexandervonhumboldt2):

11=x

OpenStudy (anonymous):

omg Thanks for helping me and giving me a medal!